A vehicle moving with a uniform acceleration of 2m/s2 has a velocity of 4m/s at a certain time.What will its velocity be? At 1s later and at 5s later


1
Expert's answer
2021-02-22T10:24:57-0500

We know that


v(t)=v0+at,v(t)=v_0+at,

so:


v(1)=4+21=6 m/s,v(5)=4+25=14 m/s.v(1)=4+2·1=6\text{ m/s},\\ v(5)=4+2·5=14\text{ m/s}.